OpenStudy (anonymous):

True-False: Please select true or false and click "submit." If t is any real number, it is possible that tan t = 5/4 .

OpenStudy (australopithecus):

tan(t) = 5/4 arctan(tan(t)) = arctan(5/4) t = arctan(5/4) does it exist?

OpenStudy (australopithecus):

type into calculator to find out

OpenStudy (anonymous):

yes

OpenStudy (australopithecus):

then you have your answer